OpenStudy (anonymous):

Graph the inequality |x| + |y| < 1 Please give the step by step solution for the above problem. Thanks in Advance

OpenStudy (phi):

for x and y >=0 (i.e. in quadrant 1) you can plot x+y=1 or in slope-intercept form: y= -x + 1 plot that line and make it dotted. be sure both x and y are positive (in other words don't plot the line so either x or y is negative) that is part of the answer

OpenStudy (phi):

then look at the case where x is positive and y is negative (quadrant 4) x -y = 1 y= x -1 plot that line in quadrant 4, and make it dotted.

OpenStudy (phi):

do the same for the other two quadrants. you will get a diamond shape, with dotted lines as the sides fill in the diamond to show all the values inside the diamond will make the original inequality true.

OpenStudy (anonymous):

Thanks for the answer